OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 3y = 10 4x + 5y = 14 2 1 -1 -2

OpenStudy (anonymous):

4x + 3y = 10 4x + 5y = 14 -------------- multiply the top equation by -1 -4x - 3y = -10 4x + 5y = 14 -------------- add vertically (-4x+4x) + (-3y+5y) = (-10+14) 0x +2y = 4 2y = 4 y = 2 put y = 2 into either equation to get the x value.